1. Two (2) small plastic spheres are given positive electrical charge.When ther 15cm apart, the repulse force between then has a magnitude of 0.22N. What is the charge on each sphere if;
  1. two (2) charges are equal
  2. one (1) sphere has four (4) times the charge of the other.

Grade:8

1 Answers

Arun
5 years ago
Just use F = kQq / d² 

a) 0.220 N = 8.99e9N·m²/C² * Q² / (0.15m)² 
Q = 7.42e-7 C = 742 nC 

b) 0.220 N = 8.99e9N·m²/C² * 4Q * Q / (0.15m)² 
Q = 3.71e-7 C 
4Q = 1.48e-6 C = 1.48 µC